# Phase Transition and Nuclear Symmetry Energy from Neutron Star Observations – In Light of PSR J0614–3329

### v1

@@ Abstract
The possible occurrence of a first-order hadron-quark phase transition (FOPT) in neutron-star interiors remains an open question. Whether such a transition can be directly tested with improved observations is a key challenge. Here, we incorporate the latest constraints, especially a new NICER radius measurement for PSR J0614–3329, into a nonparametric Gaussian Process (GP) EOS framework that explicitly includes a first-order transition. We find a Bayes factor of $B\approx 2.3$ when comparing models with and without an explicit phase transition, marginally favoring its presence. At $68\%$ credibility, the transition onset density $n_{\rm PT}$ is either below $2\,n_{s}$ (corresponding to masses $\lesssim 1\,M_{\odot}$, with density jump $\Delta n\sim 0.5\,n_{s}$) or, more prominently, above $4\,n_{s}$ (near the central density of the heaviest NS, with $\Delta n\sim 3\,n_{s}$), where $n_{s}$ represents the nuclear saturation density. In addition, by using symmetry-energy expansion at low densities ($<1.1\,n_{s}$), we infer a slope parameter $L=40.2^{+19.3}_{-14.3}$ MeV, in good agreement with nuclear-experiment values. Intriguingly, $L$ correlates positively with the radius difference between $1.4\,M_{\odot}$ and $2.0\,M_{\odot}$ stars.
@@
1 Introduction
Neutron stars (NSs) serve as natural laboratories for probing the equation of state (EOS) of cold, dense matter under extreme conditions. Many progress in constraining the EOS has been driven by multimessenger NS observations in recent years (e.g., Capano et al., 2020; Dietrich et al., 2020; Huth et al., 2022; Annala et al., 2022; Pang et al., 2023; Neill et al., 2023; Gorda et al., 2023a; Han et al., 2023; Tsang et al., 2024). The detection of gravitational waves from the binary NS merger GW170817 provided novel constraints on the tidal deformability of NSs (Abbott et al., 2018, 2019), while X-ray timing observations by NICER have yielded simultaneous mass-radius measurements for several pulsars (e.g., PSR J0030+0451 and PSR J0740+6620) (Vinciguerra et al., 2024; Salmi et al., 2024). These observations, combined with the existence of massive $\sim 2\,M_{\odot}$ pulsars (Demorest et al., 2010; Fonseca et al., 2021), have placed constraints on the stiffness of the NS EOS (Jiang et al., 2020; Landry et al., 2020; Raaijmakers et al., 2020).
Nuclear experiments and theory constrain the EOS at lower densities, especially the symmetry energy and its density dependence. Many analyses based on, e.g., nuclear masses, giant dipole resonances, and heavy‐ion collisions have converged on the symmetry energy slope $L\sim 50$ MeV (Tews et al., 2017; Drischler et al., 2020). In contrast, the PREX-II experiment reported a neutron-skin thickness in 208Pb that implies a rather large $L$, an inference that appears inconsistent with other empirical determinations (Adhikari et al., 2021; Reed et al., 2021). Meanwhile, perturbative quantum chromodynamic (pQCD) calculations at asymptotically high density provide rigorous boundary conditions that, under the requirements of thermodynamic stability and causality, can be extrapolated down toward NS‐relevant densities, generally favoring a relatively soft core in massive stars (Annala et al., 2020; Komoltsev & Kurkela, 2022; Gorda et al., 2023b; Han et al., 2023; Tang et al., 2024c; Finch et al., 2025).
An outstanding question is whether a first-order phase transition (FOPT), e.g., to deconfined quark matter, takes place in NS interiors, and if so, at what density and with what strength. Many studies has devoted considerable effort to this problem (e.g., Bauswein et al., 2019; Han & Steiner, 2019; Most et al., 2019; Montaña et al., 2019; Blacker et al., 2020; Weih et al., 2020; Miao et al., 2020; Xie & Li, 2021; Huang et al., 2022; Yue et al., 2022; Brandes et al., 2023; Essick et al., 2023; Li et al., 2023; Zhu et al., 2023; Gao et al., 2024; Prakash et al., 2024; Zhou et al., 2024; Huang & Sourav, 2025; Li et al., 2025c). A strong phase transition could lead to observable effects such as a sharp radius reduction in the NS mass-radius ($M$-$R$) relation (Lindblom, 1998; Christian et al., 2018; Blaschke et al., 2019). Therefore, it is expected the FOPT might be directly verified from a list of well-measured NS observations (Pereira et al., 2022; Mondal et al., 2023; Li et al., 2025b, d). Previous studies using parametrized EOS models and NS observations have argued against the existence of a strong FOPT at relatively low densities (Christian & Schaffner-Bielich, 2020; Tang et al., 2021a). In a model-agnostic way, Han et al. (2023) further revealed that a strong FOPT (i.e., $c_{\rm s}^{2}\rightarrow 0$) can only take place at the sufficiently-high densities of $\sim 5n_{s}$ (see their Figure 2). More recently, nonparametric approaches incorporating explicit FOPT have been applied to updated data, including chiral effective field theory ($\chi$EFT) constraints at low density (Komoltsev, 2024). These analyses suggest that a phase transition occurring at $2$-$3\,n_{s}$ is disfavored by the multimessenger NS data.
@@ 1 Introduction
2 Methods
Our EOS model builds upon our previous works: one that combined a nuclear empirical expansion with a piecewise-parametric EOS (Tang et al., 2021b), and another that implemented a Gaussian Process approach for NS matter (Tang et al., 2024b). Here we summarize the necessary details. First, in many theoretical models of cold, beta-equilibrated nuclear matter, the energy per nucleon at a given baryon density $n$ can be approximated by a quadratic expansion in isospin asymmetry:
@@ 2 Methods
where $x$ is the proton fraction. The symmetric nuclear matter (SNM) part is expanded around saturation density $n_{s}$ as
@@
[…]
Beyond $1.1\,n_{s}$, we extend the EOS using a conditioned Gaussian Process (Landry & Essick, 2019; Essick et al., 2020; Landry et al., 2020; Gorda et al., 2023b; Legred et al., 2025). Specifically, we describe the sound speed via an auxiliary variable
@@
[…]
@@
To incorporate an FOPT in this framework, we replace the GP‐derived sound‐speed profile $c_{s}^{2}(n)$ by zero over the density interval $[n_{\rm PT},\,n_{\rm PT}+\Delta n]$, thereby introducing a discontinuous jump in density. Here $n_{\rm PT}$ (the transition onset) is drawn uniformly from [1.1, $[1.1,\,10]\,n_{s}$ and $\Delta n$ (the density jump) uniformly from [0, $[0,\,8]\,n_{s}$. For each EOS sample, we impose a prior that $n_{\rm PT}$ must be less than the central density $n_{\rm TOV}$ of the maximum-mass configuration (so that any phase transition lies within the stable branch of NSs). We also discard any EOS sample where $n_{\rm PT}+\Delta n>10\,n_{s}$. Additionally, we constrain the maximum nonrotating NS mass $M_{\rm TOV}$ to lie between $2\,M_{\odot}$ and $3\,M_{\odot}$.
@@
3 Results
Figure 1: Prior (grey) and posterior (blue) joint distributions of the phase transition onset density $n_{\rm PT}$ and the density jump $\Delta n$.
Figure 2: Mass-radius distributions for NSs whose central densities correspond to the low (orange) and high (blue) modes of the phase‐transition onset density in Figure 1. The black dashed curve shows the observed Galactic NS mass distribution, and the vertical dot-dash line marks the lowest reliably measured mass, that of the companion to PSR J0453+1559.
Figure 3: Prior (grey) and posterior (blue) joint distributions of $n_{\rm PT}$ and the radius of the maximum-mass NS ($R_{\rm TOV}$).
Figure 4: Prior (grey) and posterior (blue) joint distributions of the symmetry energy slope $L$ and the radius difference $R_{1.4}-R_{2.0}$.
@@ 3 Results
Since a strong FOPT might be directly observable (e.g, a sharp radius reduction), we examine the NS masses at which the transition starts. Figure 2 shows the mass-radius distributions for neutron stars whose central densities coincide with the two posterior modes of the phase‐transition onset: the low‐density mode ($n_{\rm PT}<2.5\,n_{s}$, orange) and the high‐density mode (blue). The low‐density branch falls below $\sim 1\,M_{\odot}$ (so lower than the lightest NS PSR J0453+1559 (Martinez et al., 2015)), while the high‐density branch lies near the maximum‐mass configuration. Both regimes lie well outside the bulk of the observed Galactic neutron‐star mass distribution (black dashed curve, Shao et al., 2020; Fan et al., 2024), indicating that stars which sample these phase‐transition densities are exceedingly rare. Therefore, standard mass-radius measurements of typical pulsars are unlikely to directly probe the phase‐transition parameter space favored by our analysis.
Intriguingly, we find a positive correlation between the phase transition onset density and the radius of the maximum-mass NS ($R_{\rm TOV}$). As shown in Figure 3, EOS samples with higher $n_{\rm PT}$ tend to yield larger values of $R_{\rm TOV}$. In our earlier work, we suggested that the radius of the most massive NS could be essential for constraining high-density EOS (Tang et al., 2024b). The present results reinforce this idea: if $R_{\rm TOV}$ could be determined in a good accuracy, for instance, via post-merger gravitational-wave signals (Breschi et al., 2019) or supramassive NSs originated black holes (Tang et al., 2024a), it would help pinpoint the density at which a possible transition occurs.
@@
Figure 5: The $68.3\%$ credible ranges for the pressure (as a function of $n$; top) and the squared sound speed $c_{s}^{2}$ (bottom), with (blue) and without (grey) the inclusion of PSR J0614–3329. The orange lines show results from EOS models without an explicit FOPT.
Figure 6: The $68.3\%$ credible neutron star mass-radius intervals, compared to observational radius measurements. Blue curves and bands denote results including PSR J0614–3329, while grey curves correspond to results without J0614. The orange lines show results from EOS models without an explicit FOPT.
Figures 5 and 6 illustrate the impact of incorporating the new PSR J0614–3329 measurement on the EOS and $M$-$R$ posteriors. In Figure 5, we show the $68.3\%$ credible bands for the pressure as a function of density and for the corresponding sound speed $c_{s}^{2}(n)$. The inclusion of J0614 (blue bands) leads to a slightly softer EOS around $2\,n_{s}$, compared to the previous constraints (grey bands). In particular, the pressure at intermediate densities now increases more gradually (whereas the previous analysis favored a somewhat more rapid stiffening around $2$-$3\,n_{s}$). The behavior of $c_{s}^{2}(n)$ still exhibits a pronounced peak, and the inferred phase transition tends to occur at relatively high densities, consistent with the $n_{\rm PT}$ posterior.
@@
Figure 6 shows the reconstructed NS mass-radius relation ($68.3\%$ credible region) alongside measured NS radius data sets. We find that adding the PSR J0614–3329 constraint shifts the overall radius estimates toward smaller values (blue region) by roughly a few hundred meters in comparison to the result without J0614 (grey region). Notably, the updated median radius for a $1.4\,M_{\odot}$ NS is now in closer alignment with the lower-radius mode inferred from GW170817 analyses. We also observe that the high-mass end of the $M$-$R$ curve tends to bend toward slightly larger radii (visible as a subtle upturn of the blue band for $M\gtrsim 2.25\,M_{\odot}$).
@@
Figure 7: Prior (grey) and posterior (blue) distributions of selected EOS parameters.
@@
Figure 8: Prior (grey) and posterior (blue) distributions of neutron star bulk properties.
@@
In Figure 7, we present the prior vs. posterior distributions for several EOS parameters. In addition to the constraints on $n_{\rm PT}$ and $\Delta n$ discussed above, we find that the symmetry energy slope $L$ is tightly constrained by the data to $40.18_{-14.30}^{+19.25}$ MeV, which is smaller than (though still consistent with) that summarized in Li et al. (2021). The posterior probability for higher $L$ values is significantly suppressed compared to the broad prior. Likewise, the symmetry energy curvature $K_{\rm sym}$ is pulled toward negative values, with only a small probability for positive $K_{\rm sym}$ remaining. Figure 8 shows the distributions for key NS bulk properties. The maximum nonrotating mass $M_{\rm TOV}$ is inferred to be $\sim 2.25\,M_{\odot}$, and the corresponding radius is about $R_{\rm TOV}\approx 11.6$ km, confirming our previous conclusions (Fan et al., 2024; Tang et al., 2025) and indicating the robustness of this constraint. The radii of canonical-mass NSs (e.g., $R_{1.4}$ and $R_{2.0}$) are also well constrained, with uncertainties on the order of only a few hundred meters. Notably, the posterior for the radius difference $\Delta R=R_{1.4}-R_{2.0}$ is centered at a negative value (indicating $R_{2.0}>R_{1.4}$), whereas the prior allowed a wider range including large positive $\Delta R$ values that are now strongly disfavored. A negative $\Delta R$ is a potential indicator of a phase transition, as it reflects an EOS that becomes softer at higher densities (causing the most massive stars to have relatively smaller radii). Finally, the tidal deformability of a $1.4\,M_{\odot}$ NS, $\Lambda_{1.4}$, is constrained to be on the order of $\sim 300$, in line with the tighter radius constraints.
4 Summary

### v2

# Phase transition and nuclear symmetry energy from neutron star observations: Constraints in light of PSR J0614–3329

@@ Abstract
The possible occurrence of a first-order hadron-quark phase transition in neutron-star interiors remains an open question. Whether such a transition can be directly tested with improved observations is a key challenge. Here, we incorporate the latest constraints, especially a new NICER radius measurement for PSR J0614–3329, into a nonparametric Gaussian process equation of state framework that explicitly includes a first-order transition. We find a Bayes factor of $B\approx 2.3$ when comparing models with and without an explicit phase transition, marginally favoring its presence. At $68\%$ credibility, the transition onset density $n_{\rm PT}$ is either below $2\,n_{s}$ (corresponding to masses $\lesssim 1\,M_{\odot}$, with density jump $\Delta n\sim 0.5\,n_{s}$) or, more prominently, above $4\,n_{s}$ (near the central density of the heaviest neutron star, with $\Delta n\sim 3\,n_{s}$), where $n_{s}$ represents the nuclear saturation density. In addition, by using symmetry-energy expansion at low densities ($1.1\,n_{s}$), we infer a slope parameter $L=40.2^{+19.3}_{-14.3}$ MeV, in good agreement with nuclear-experiment values. Intriguingly, $L$ correlates positively with the radius difference between $1.4\,M_{\odot}$ and $2.0\,M_{\odot}$ stars.
I Introduction
Neutron stars (NSs) serve as natural laboratories for probing the equation of state (EOS) of cold, dense matter under extreme conditions. Much progress in constraining the EOS has been driven by multimessenger NS observations in recent years, e.g., Refs. (Capano et al., 2020; Dietrich et al., 2020; Huth et al., 2022; Annala et al., 2022; Pang et al., 2023; Neill et al., 2023; Gorda et al., 2023a; Han et al., 2023; Tsang et al., 2024). The detection of gravitational waves from the binary NS merger GW170817 provided novel constraints on the tidal deformability of NSs (Abbott et al., 2018, 2019), while x-ray timing observations by the Neutron star Interior Composition Explorer (NICER) have yielded simultaneous mass-radius measurements for several pulsars (e.g., PSR J0030+0451 and PSR J0740+6620) (Vinciguerra et al., 2024; Salmi et al., 2024). These observations, combined with the existence of massive $\sim 2\,M_{\odot}$ pulsars (Demorest et al., 2010; Fonseca et al., 2021), have placed constraints on the stiffness of the NS EOS (Jiang et al., 2020; Landry et al., 2020; Raaijmakers et al., 2020).
Nuclear experiments and theory constrain the EOS at lower densities, especially the symmetry energy and its density dependence. Many analyses based on, e.g., nuclear masses, giant dipole resonances, and heavy‐ion collisions have converged on the symmetry energy slope $L\sim 50$ MeV (Tews et al., 2017; Drischler et al., 2020). In contrast, the Lead Radius Experiment-II (PREX-II) reported a neutron-skin thickness in 208Pb that implies a rather large $L$, an inference that appears inconsistent with other empirical determinations (Adhikari et al., 2021; Reed et al., 2021). Meanwhile, perturbative quantum chromodynamic (pQCD) calculations at asymptotically high density provide rigorous boundary conditions that, under the requirements of thermodynamic stability and causality, can be extrapolated down toward NS‐relevant densities, generally favoring a relatively soft core in massive stars (Annala et al., 2020; Komoltsev and Kurkela, 2022; Gorda et al., 2023b; Han et al., 2023; Tang et al., 2024a; Finch et al., 2025).
@@ I Introduction
II Methods
Our EOS model builds upon our previous works: one that combined a nuclear empirical expansion with a piecewise-parametric EOS (Tang et al., 2021b), and another that implemented a Gaussian process approach for NS matter (Tang et al., 2024b). Here we summarize the necessary details. First, in many theoretical models of cold, beta-equilibrated nuclear matter, the energy per nucleon at a given baryon density $n$ can be approximated by a quadratic expansion in isospin asymmetry, as follows:
@@ II Methods
where $x$ is the proton fraction. The symmetric nuclear matter part is expanded around saturation density $n_{s}$ as
@@
where $m_{\rm p,n}$ are the proton and neutron rest masses, $\varepsilon_{e}$ is the energy density of a degenerate relativistic electron gas, and the proton fraction $x$ is determined by minimizing the total energy at a given $n$. We then match this EOS to the crust model at lower densities (Baym et al., 1971; Negele and Vautherin, 1973). The crust-core transition density is identified by the vanishing of the effective incompressibility of npe($\mu$) matter at beta equilibrium and charge neutrality conditions (Kubis, 2007; Lattimer and Prakash, 2007; Zhang et al., 2018). In our Bayesian framework, we sample the incompressibility coefficient $K_{0}$ from a normal distribution $\mathcal{N}(240,\,30^{2})$ truncated to the interval [210, 270][210,\,270] MeV, the symmetry energy $S_{v}$ at saturation from $\mathcal{N}(31.7,\,3.2^{2})$ truncated to [22, 44][22,\,44] MeV, the slope parameter $L$ from a uniform distribution on [20, 120][20,\,120] MeV, and the symmetry‐energy curvature $K_{\rm sym}$ from a uniform distribution on [400, 100][-400,\,100] MeV, while fixing the binding energy at saturation to $E_{0}(n_{s})=-15.9$ MeV.
Beyond $1.1\,n_{s}$, we extend the EOS using a conditioned Gaussian process (Landry and Essick, 2019; Essick et al., 2020; Landry et al., 2020; Gorda et al., 2023b; Legred et al., 2025). Specifically, we describe the sound speed via an auxiliary variable
@@
which is treated as a multivariate Gaussian distribution. That is, $\phi(n)\sim\mathcal{N}\big(-\ln(1/\bar{c}_{s}^{2}-1),\,K(n,n^{\prime})\big)$, with a Gaussian kernel $K(n,n^{\prime})=\eta\exp[-(n-n^{\prime})^{2}/(2l^{2})]$. The three hyperparameters of this GP, namely the variance $\eta$, correlation length $l$, and mean sound-speed-squared $\bar{c}_{s}^{2}$, are drawn from hyperprior distributions: $\eta\sim\mathcal{N}(1.25,\,0.2^{2})$, $l\sim\mathcal{N}(0.5\,n_{s},\,(0.25\,n_{s})^{2})$, and $\bar{c}_{s}^{2}\sim\mathcal{N}(0.5,\,0.25^{2})$. We condition the GP on the low-density EOS segment as follows: let $\phi_{\rm nuc}$ denote the values of $\phi(n)$ computed from the nuclear expansion EOS at densities $n_{\rm nuc}\leq 1.1\,n_{s}$. We treat these as “data” points for the GP with an assumed variance $\sigma^{2}_{\phi_{\rm nuc}}=10^{-4}$ at each $n_{\rm nuc}$. The conditioned GP then yields
@@
To incorporate an FOPT in this framework, we replace the GP‐derived sound‐speed profile $c_{s}^{2}(n)$ by zero over the density interval $n_{\rm PT},\,n_{\rm PT}+\Delta n]$, thereby introducing a discontinuous jump in density. This corresponds to a Maxwell-type FOPT (sharp interface, with $c_{s}^{2}\to 0$). We do not consider the alternative Gibbs-type first-order transition, which would involve a mixed phase and a nonvanishing sound speed. Here $n_{\rm PT}$ (the transition onset) is drawn uniformly from [1.1, $1,\,10]\,n_{s}$ and $\Delta n$ (the density jump) uniformly from $0,\,8]\,n_{s}$. For each EOS sample, we impose a prior that $n_{\rm PT}$ must be less than the central density $n_{\rm TOV}$ of the maximum-mass configuration (so that any phase transition lies within the stable branch of NSs). We also discard any EOS sample where $n_{\rm PT}+\Delta n>10\,n_{s}$. Additionally, we constrain the maximum nonrotating NS mass $M_{\rm TOV}$ to lie between $2\,M_{\odot}$ and $3\,M_{\odot}$.
@@
III Results
Figure 1: Prior (gray) and posterior (blue) joint distributions of the phase transition onset density $n_{\rm PT}$ and the density jump $\Delta n$.
Figure 2: Mass-radius distributions for NSs whose central densities correspond to the low (orange) and high (blue) modes of the phase‐transition onset density in Fig. 1. The black dashed curve shows the observed Galactic NS mass distribution, and the vertical dot-dash line marks the lowest reliably measured mass, that of the companion to PSR J0453+1559.
Figure 3: Prior (gray) and posterior (blue) joint distributions of $n_{\rm PT}$ and the radius of the maximum-mass NS ($R_{\rm TOV}$).
Figure 4: Prior (gray) and posterior (blue) joint distributions of the symmetry energy slope $L$ and the radius difference $R_{1.4}-R_{2.0}$.
@@ III Results
Since a strong FOPT might be directly observable (e.g, a sharp radius reduction), we examine the NS masses at which the transition starts. Figure 2 shows the mass-radius distributions for neutron stars whose central densities coincide with the two posterior modes of the phase‐transition onset: the low‐density mode ($n_{\rm PT}<2.5\,n_{s}$, orange) and the high‐density mode (blue). The low‐density branch falls below $\sim 1\,M_{\odot}$ (so lower than the lightest NS PSR J0453+1559 (Martinez et al., 2015)), while the high‐density branch lies near the maximum‐mass configuration. Both regimes lie well outside the bulk of the observed Galactic neutron‐star mass distribution (black dashed curve, (Shao et al., 2020; Fan et al., 2024)), indicating that stars which sample these phase‐transition densities are exceedingly rare. Therefore, standard mass-radius measurements of typical pulsars are unlikely to directly probe the phase‐transition parameter space favored by our analysis.
Intriguingly, we find a positive correlation between the phase transition onset density and the radius of the maximum-mass NS ($R_{\rm TOV}$). As shown in Fig. 3, EOS samples with higher $n_{\rm PT}$ tend to yield larger values of $R_{\rm TOV}$. In our earlier work, we suggested that the radius of the most massive NS could be essential for constraining high-density EOS (Tang et al., 2024b). The present results reinforce this idea: if $R_{\rm TOV}$ could be determined in a good accuracy, for instance, via postmerger gravitational-wave signals (Breschi et al., 2019) or supramassive NS-originated black holes (Tang et al., 2024c), it would help pinpoint the density at which a possible transition occurs.
@@
Figure 5: $68.3\%$ credible ranges for the pressure (as a function of $n$; top) and the squared sound speed $c_{s}^{2}$ (bottom), with (blue) and without (gray) the inclusion of PSR J0614–3329. The orange lines show results from EOS models without an explicit FOPT.
Figure 6: $68.3\%$ credible neutron star mass-radius intervals, compared to observational radius measurements. Blue curves and bands denote results including PSR J0614–3329, while gray curves correspond to results without PSR J0614–3329. The orange lines show results from EOS models without an explicit FOPT.
Figures 5 and 6 illustrate the impact of incorporating the new PSR J0614–3329 measurement on the EOS and $M$-$R$ posteriors. In Fig. 5, we show the $68.3\%$ credible bands for the pressure as a function of density and for the corresponding sound speed $c_{s}^{2}(n)$. The inclusion of J0614 (blue bands) leads to a slightly softer EOS around $2\,n_{s}$, compared to the previous constraints (gray bands). In particular, the pressure at intermediate densities now increases more gradually (whereas the previous analysis favored a somewhat more rapid stiffening around 22-$3\,n_{s}$). The behavior of $c_{s}^{2}(n)$ still exhibits a pronounced peak, and the inferred phase transition tends to occur at relatively high densities, consistent with the $n_{\rm PT}$ posterior.
@@
Figure 6 shows the reconstructed NS mass-radius relation ($68.3\%$ credible region) alongside measured NS radius datasets. We find that adding the PSR J0614–3329 constraint shifts the overall radius estimates toward smaller values (blue region) by roughly a few hundred meters in comparison to the result without J0614 (gray region). Notably, the updated median radius for a $1.4\,M_{\odot}$ NS is now in closer alignment with the lower-radius mode inferred from GW170817 analyses. We also observe that the high-mass end of the $M$-$R$ curve tends to bend toward slightly larger radii (visible as a subtle upturn of the blue band for $M\gtrsim 2.25\,M_{\odot}$).
@@
Figure 7: Prior (gray) and posterior (blue) distributions of selected EOS parameters.
Figure 8: Prior (gray) and posterior (blue) distributions of EOS parameters when the symmetry-energy expansion is extended to $1.85\,n_{s}$.
@@
Figure 9: Prior (gray) and posterior (blue) distributions of neutron star bulk properties.
[…]
@@
Figure 9 shows the distributions for key NS bulk properties. The maximum nonrotating mass $M_{\rm TOV}$ is inferred to be $\sim 2.25\,M_{\odot}$, and the corresponding radius is about $R_{\rm TOV}\approx 11.6$ km, confirming our previous conclusions (Fan et al., 2024; Tang et al., 2025) and indicating the robustness of this constraint. The radii of canonical-mass NSs (e.g., $R_{1.4}$ and $R_{2.0}$) are also well constrained, with uncertainties on the order of only a few hundred meters. Notably, the posterior for the radius difference $\Delta R=R_{1.4}-R_{2.0}$ is centered at a negative value (indicating $R_{2.0}>R_{1.4}$), whereas the prior allowed a wider range including large positive $\Delta R$ values that are now strongly disfavored. A negative $\Delta R$ is a potential indicator of a phase transition, as it reflects an EOS that becomes softer at higher densities (causing the most massive stars to have relatively smaller radii). Finally, the tidal deformability of a $1.4\,M_{\odot}$ NS, $\Lambda_{1.4}$, is constrained to be on the order of $\sim 300$, in line with the tighter radius constraints.
IV Summary